# Vulnerability of Text-Matching in ML/AI Conference Reviewer Assignments to Collusions

### v1

@@ 5.7 Correlation of Attack Success between NeurIPS 2022 and NeurIPS 2023
Figure 7: Manipulated rankings in 2022 and 2023 iterations of NeurIPS are strongly correlated, so colluders can estimate the manipulated rankings using previous year’s data. Spearman’s rank correlation coefficients $\rho$ are reported, and the dotted lines are $y=x$.
An attacker can use publicly available data from the previous year’s conference to train or validate their attack. For example, in Section 5.2, we discussed an early stopping heuristic that assesses the effectiveness of our attack. This involves iteratively modifying the abstract, and halting the modifications if the colluding reviewer becomes the most similar reviewer to the colluding paper among all reviewers from the prior year. The relationship between success of any attack in the previous year to the success in the target year is not clear a priori, and we investigate it in this section.
In more detail, we calculate the manipulated rankings amongst both NeurIPS 2022 and NeurIPS 2023 reviewer pools for all colluding (paper $p$, reviewer $r$) pairs from Section 5.3. We plot the attack successes in Figure 7 for three values of the natural ranking – 101, 501 and 1001. We discover a strong correlation between the manipulated rankings in the 2022 and 2023 iterations of NeurIPS. The strong correlation of attack success between NeurIPS 2022 (publicly available) and NeurIPS 2023 (unknown to attackers) reviewer pools implies that adversarial attackers can estimate attack success using previous year’s data. The reviewer pools data many major ML/AI conferences publish gives attackers a dataset to carefully tune their modifications, knowing that being successful on the public data is often good enough.
@@ 5.8 Human Subject Experiment on Detectability
In the previous sections, we found that our method can successfully increase the manipulated rankings between a paper and a colluding reviewer. However, the paper may also be assigned to honest reviewers, and in this section we describe a randomized control trial that we conduct to understand the perception of adversarial abstracts by unsuspecting human reviewers. Our research question and study design were pre-registered at https://aspredicted.org/HXF_Z92. This experiment was approved by an independent Institutional Review Board at Carnegie Mellon University (Federal Wide Assurance No: FWA00004206, IRB Registration No: IRB00000603). The broad research question we looked to answer was: Are benign (control) and adversarial (experimental) abstracts the same to unsuspecting human reviewers? We also acknowledge that, as discussed below, this experimental design also had several limitations that can lead to an overestimation of the detectability of the adversarially modified abstract.
@@ 5.8.1 Experiment design
We recruited 59 participants with one opting out after debriefing (we initially employ deception to hide the true purpose of this study), making it a total of 58 participants. The participants are all either PhD students or those who have graduated with a PhD – who have reviewed at ML/AI conferences in the past five years. Participants were recruited by emails and word of mouth at several U.S. universities, and each participant was compensated \$20 for their participation. To simulate the behavior of an unsuspecting reviewer at an actual conference, we employ deception so that the participants have no knowledge of the adversarial manipulations prior to starting the study.
We curated two pools of papers — benign and adversarial, each containing modified versions of 914 real machine learning papers we collected from arxiv.org. To generate the adversarial version of each paper, we first select a reviewer that has natural ranking of 101 out of the NeurIPS 2023 reviewer pool, then we curate the adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$ (Section 4.1) and generate $a^{\text{adv}}_{p}$ by Algorithm 1 and Algorithm 2 under the fully automatic mode for scalability. As a result of the adversarial attack, the colluding reviewer for each arXiv paper with an adversarial abstract had a manipulated ranking within the top-5 of all NeurIPS 2023 reviewers. Since the adversarial abstracts are completely LLM-generated without any human oversight, we believe that directly comparing them against the original human-written abstracts may introduce a confounder to our experiment that the adversarial abstracts are LLM-generated while the benign abstracts may be written by humans. To address this, in the benign (control) condition, we ask the LLM to paraphrase the original abstract.
Each participant was assigned two personalized papers based on their areas of expertise. Each paper assignment was chosen to be in either the benign (control) or adversarial (experimental) condition uniformly at random. Depending on whether the assignment is control or experimental, we computed the participant’s SPECTER similarities to the pool of benign or adversarial papers respectively. Using this process, we assigned each participant two distinct papers they are the most similar to.
@@ 5.8.2 Attack budgets and LLM prompts
For the human study experiment, we chose $\bm{N}=3$, ${\bm{M}}=2$ and ${\bm{K}}=5$, and the prompts can be found in Appendix B.3. Firstly, in the IncludeThemes operation, we prompt the LLM to follow the same rules and add only one sentence about the colluding reviewer’s archive, just like in the rest of this paper. In the InsertKeywords operation, we ask the LLM to add 12 keywords in this experiment, which is more than the 10 keywords ($\bm{N}=2$, ${\bm{K}}=5$) we use in the rest of this paper. Finally, early stopping is also not used in the automatic abstracts modification process for this experiment. Due to the extra keywords added, the adversarially modified abstract tested in this experiment may be more detectable than it would be for the abstracts generated in previous sections.
@@ 5.8.5 Limitations
  • We generated the adversarial abstracts automatically without any degree of human oversight for scalability. In the real-world scenario, malicious authors can at least look over the manipulated abstract to check for detectability.
@@ 5.8.6 Results
We collected a total of 116 mini-reviews from participants, comprising 49 reviews for the control group and 67 reviews for the experimental group. Out of 116 mini-reviews, 51 of the reviews included free response comments for Question 4 in the survey (Section 5.8.4). As we discuss below, the results are mixed. First, we consider the total counts of collected Yes/No responses about the paper-reviewer expertise alignment as well as the consistency and coherence of the abstracts. Table 5 summarizes the findings, where we see that control abstracts are considered significantly more consistent than experimental abstracts, but no significant difference is found in expertise alignment and abstract coherence.
@@
In Table 6 we analyze the text comments provided. We identified five main types of complaints in the comments, and report in Table 6 the percentages of reviews in control and experimental groups that mention each type of complaint. The percentages are out of all reviews collected for each group (49 reviews for control and 67 reviews for experimental), and there may be multiple types of complaints mentioned in each comment. The most prevalent issue for both benign and adversarial abstracts is writing style, which is related to word choice, tone, and readability. Surprisingly, complaints like “contains things never mentioned in the paper” exist for both benign and adversarial abstracts. To investigate whether these complaints are due to LLM hallucinations during the benign paraphrasing of control abstracts or directly due to the original contents uploaded to arXiv, we manually examined the paraphrased control abstract and the original abstract related to each complaint. We find that all types of complaints have at least one paper which we believe the complaint applies to the original abstract. Lastly, we investigated whether the review comments for the adversarial abstracts mentioned things directly related to the modification operations proposed in this work. In the experimental group, we noticed that 6% of the responses mentioned irregularities related to IncludeThemes, 16% of the responses mentioned irregularities related to InsertKeywords, and 82% of the responses are not directly related to either operations. Some responses may mention both operations, so the percentages do not sum to 100%. Finally, none of the participants identified malicious intent in any of the abstracts.
@@ 6 Discussion
In this work, we identify the vulnerability of current text-based reviewer-assignment systems in AI/ML venues to manipulation by collusion rings. This contradicts common perceptions that text-based automatic reviewer matching is resistant to such tactics. Our work suggests that a simple and practical attack procedure can effectively manipulate paper-reviewer similarities and hence manipulate automated reviewer assignment at many ML/AI conferences. Our attacks have a high success rate in matching to a targeted colluding reviewer. In the human subject experiment testing for attack detectability, no participant detected malicious intent for the manipulated abstracts. While there were complaints about coherence and consistency, benign abstracts edited by LLMs elicited similar complaints, suggesting plausible deniability. In total, we underscore the need for enhanced robustness in reviewer assignment algorithms to protect the integrity of the peer review process.
@@ 6.1 Recommendations for mitigation
  • Increase reviewer profile requirements: Our results indicate that requiring reviewers to have more extensive publication archives reduces the effectiveness of manipulation. This approach makes it more difficult for colluding reviewers to align their profiles strategically with targeted abstracts.
@@
  • Increase reviewer awareness: Educating reviewers about the possibility of such manipulations can prompt them to be more vigilant when evaluating abstracts and papers.
  • Introduce randomness in assignments: No component of automated assignment methods is entirely immune to manipulation. This highlights the importance of adopting a broader mitigation strategy such as that proposed in [JZL+20] of introducing a degree of randomness into the assignment process. Such randomness provably prevents adversaries from reliably influencing assignments.
  • Consider robustness of similarity scores Developers of similarity computation algorithms should pay attention to impart some robustness so the similarity scores are less susceptible to adversarial manipulations to the abstract.
@@ 6.2 Limitations
While our study provides valuable insights, it has several limitations that offer avenues for future research. Our analysis considers scenarios where authors collude with a single reviewer within a panel of 3-6 reviewers. Expanding this to cases where multiple reviewers colluding with the author simultaneously could provide a more comprehensive understanding of the system’s vulnerabilities. We also did not factor in common constraints such as reviewer and paper load limits, conflicts of interest, or geographical considerations. Incorporating these constraints could affect the attack’s efficacy and the generalizability of some of our findings. Finally, it would be useful to evaluate the attack success rates in other venues.
In our human subject experiment, the adversarially modified abstract tested may be more detectable than it would be for the adversarial abstracts evaluated in all other experiments, due to the higher attack budget that was erroneously tuned for the human subject experiment. In addition, in a laboratory experiment like ours, we could not fully replicate the complexities of real-world reviewer behavior when presented with adversarially modified abstracts. Factors such as varying levels of expertise, bias, and attention could influence outcomes in practical settings. We also do not consider how different reviewers could influence each other.
6.3 Ethical Considerations
The primary purpose of this study is to better understand and illuminate the weaknesses in existing reviewer matching systems, not to facilitate or encourage malicious attacks. To promote transparency and foster improvements in the system, we have made our attack algorithm and adversarial abstract examples publicly available. We have also informed the concerned conference management platform about these results. Collusions and attacks as described in this paper could already be taking place in practice, but most researchers may not be aware of these vulnerabilities. We also do not know of any actual investigations on detecting any such existing behavior. Therefore, we believe the benefits of this research outweigh the risk. By raising awareness of these vulnerabilities, we aim to empower non-colluding reviewers to be more vigilant when assessing abstracts. Enhanced awareness can lead reviewers to scrutinize submissions more carefully, potentially identifying and addressing any unwarranted influences.
We recognize that such attacks often come with plausible deniability. It is challenging to prove intentional malfeasance on the part of authors or reviewers based solely on abstract content or profile changes. This inherent ambiguity underscores the importance of proactively identifying potential sources of such attacks and mitigating the feasibility of such attacks, rather than relying solely on post hoc detection and enforcement.
All in all, we hope that our findings and recommendations will help the research community be more vigilant and robust to such malicious activities in scientific peer review.
@@ Appendix A Adversarial Abstract Modification Algorithms
    SimilarityCheck($t_{p}$,${a^{\text{adv}}_{p}}^{\prime}$,$a^{\text{adv}}_{p}$,$\mathcal{Q}_{r}^{\text{adv}}$,$\delta$) queries the SPECTER model and returns true if the similarity of abstract ${a^{\text{adv}}_{p}}^{\prime}$ is higher than (or at least comparable to) the similarity of abstract $a^{\text{adv}}_{p}$, that is, $s((t_{p},{a^{\text{adv}}_{p}}^{\prime}),\mathcal{Q}_{r}^{\text{adv}})+\delta>s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$. The $\delta$ parameter represents a small non-negative value, for cases when the new edits added in ${a^{\text{adv}}_{p}}^{\prime}$ do not have to be more similar to reviewer $r$ than the $a^{\text{adv}}_{p}$.
@@
    EarlyStoppingCheck($a^{\text{adv}}_{p}$,$\mathcal{Q}_{r}^{\text{adv}}$) returns true if the colluding reviewer $r$ (with the curated adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$) is the most-similar reviewer for the paper $p$ (with the adversarial abstract $a^{\text{adv}}_{p}$) amongst some (potentially proxy) set of reviewers. In the proposed method, abstracts are modified in a multistep process, so an early stopping check may be desirable to stop further modifications if the attack is envisaged to be successful. The use of early stopping is optional in our algorithm, since it is designed to trade off between attack effectiveness and abstract modification strength. In one of our experiments, we use early stopping with a proxy set of reviewers from the previous iteration of the conference: if the colluding reviewer is the most-similar reviewer for the colluding paper among all reviewers in the previous edition of the conference, no further abstract modifications are made. This heuristic is feasible and realistic because many major AI/ML conferences publish their reviewer pools from previous years. We validate the effectiveness of our early stopping heuristic in Section 5.2 and further investigate the use of proxy sets in Section 5.7.

### v2

@@ 5.6 Correlation of Attack Success between NeurIPS 2022 and NeurIPS 2023
Figure 5: Manipulated rankings in 2022 and 2023 iterations of NeurIPS are strongly correlated, so colluders can estimate the manipulated rankings using previous year’s data. Spearman’s rank correlation coefficients $\rho$ is reported, and the dotted line is $y=x$.
An attacker can use publicly available data from the previous year’s conference to train or validate their attack. For example, in Section 5.2, we discussed an effective early stopping heuristic that halts further modifications once the colluding reviewer becomes the top match amongst prior-year reviewers. The relationship between success of any attack in the previous year to the success in the target year is not clear a priori, and we investigate it in this section.
In Figure 5, we calculate the manipulated rankings amongst both NeurIPS 2022 and NeurIPS 2023 reviewer pools for 300 colluding (paper $p$, reviewer $r$) pairs with natural rankings of 101 and 100 $(p,r)$ pairs each for natural rankings of 20, 501, and 1001 from Section 5.3. As shown in Figure 5, we discover a strong correlation between the manipulated rankings in the 2022 (publicly available) and 2023 (unknown to attackers) iterations of NeurIPS. This implies that adversarial attackers can estimate attack success using previous year’s data. The reviewer pools data many major ML/AI conferences publish give attackers a dataset to tune their modifications, knowing that being successful on the public data is often good enough.
@@ 5.7 Human Subject Experiment on Detectability
In the previous sections, we found that our method can successfully increase the manipulated rankings between a paper and a colluding reviewer. However, the paper may also be assigned to honest reviewers, and in this section we describe a randomized control trial that we conduct to understand the perception of adversarial abstracts by unsuspecting human reviewers. The research question and study design were preregistered at https://aspredicted.org/HXF_Z92. This experiment was approved by an independent Institutional Review Board at Carnegie Mellon University (Federal Wide Assurance No: FWA00004206, IRB Registration No: IRB00000603). The broad research question we looked to answer was: Are benign (control) and adversarial (experimental) abstracts the same to unsuspecting human reviewers? We also acknowledge that, as discussed below, this experimental design also had several limitations that can lead to an overestimation of the detectability of the adversarially modified abstract.
@@ 5.7.1 Experiment design
We recruited 59 participants with one opting out after debriefing (we initially employ deception to hide the true purpose of this study), making it a total of 58 participants. The participants are all either PhD students or those who have graduated with a PhD – who have reviewed at ML/AI conferences in the past five years. Participants were recruited by emails and word of mouth at several U.S. universities, and each participant was compensated \$20. To simulate the behavior of an unsuspecting reviewer at an actual conference, we employ deception so that the participants have no knowledge of the adversarial manipulations prior to starting the study.
We curated two pools of papers — benign and adversarial, each containing modified versions of 914 real machine learning papers we collected from arxiv.org. To generate the adversarial version of each paper, we first select a reviewer that has natural ranking of 101 out of the NeurIPS 2023 reviewer pool, then we curate the adversarial archive $Q_{r}^{\text{adv}}$ (Section 4.1) and generate $a^{\text{adv}}_{p}$ by Algorithm 1 and Algorithm 2 under the fully automatic mode for scalability. As a result of the adversarial attack, the colluding reviewer for each arXiv paper with an adversarial abstract had a manipulated ranking within the top-5 of all NeurIPS 2023 reviewers. Since the adversarial abstracts are LLM-generated without any human oversight, we believe that directly comparing them against the original human-written abstracts may introduce a confounder to our experiment that the adversarial abstracts are LLM-generated while the benign abstracts may be written by humans. To address this, in the benign (control) paper pool, we ask the LLM to paraphrase the original abstract.
Each participant was assigned two personalized papers based on their expertise. Each paper assignment was chosen to be in either the benign (control) or adversarial (experimental) condition uniformly at random. Depending on whether the assignment is control or experimental, we computed the participant’s SPECTER similarities to the pool of benign or adversarial papers, respectively. Using this process, we assigned each participant two distinct papers they are the most similar to.
@@ 5.7.2 Attack budgets and LLM prompts
For the human study experiment, we chose $\bm{N}=3$, ${\bm{M}}=2$ and ${\bm{K}}=5$, and the prompts can be found in Appendix E.3. Firstly, in the IncludeThemes operation, we prompt the LLM to follow the same rules and add only one sentence about the colluding reviewer’s archive, just like in the rest of this paper. In the InsertKeywords operation, we ask the LLM to add 12 keywords in this experiment, which is more than the 10 keywords ($\bm{N}=2$, ${\bm{K}}=5$) we use in the rest of this paper. Finally, early stopping is also not used in the automatic abstracts modification process for this experiment. Due to the extra keywords added, the adversarially modified abstract tested in this experiment may be more detectable than it would be for the abstracts generated in previous sections.
@@ 5.7.5 Limitations
  • We generated the adversarial abstracts automatically without any degree of human oversight for scalability. This may not reflect operation of actual colluders in the real world, where malicious authors can at least look over the manipulated abstract to check for detectability.
@@
Finally, we note that the monetary compensation may have induced some participants to pay less attention than others (e.g., if their participation was solely for compensation).
@@ 5.7.6 Results
We collected a total of 116 mini-reviews from participants, comprising 49 reviews for the control group and 67 reviews for the experimental group. Out of 116 mini-reviews, 51 of the reviews included free response comments for Question 4 in the survey (Section 5.7.4). As we discuss below, the results are mixed. First, we consider the total counts of Yes/No responses about the paper-reviewer expertise alignment as well as the consistency and coherence of the abstracts. Table 5 summarizes the findings, where we see that control abstracts are considered significantly more consistent than experimental abstracts, but no significant difference is found in expertise alignment and abstract coherence.
@@
In Table 6 we analyze the free response comments provided. Each review that indicated either poor coherence or inconsistency includes comments, and all comments are between 7 and 189 words. We identify five main types of complaints in the comments, and we report the percentages of reviews in control and experimental groups that mention each type of complaint. The percentages are out of all reviews collected for each group (49 control reviews and 67 experimental reviews), and there may be multiple types of complaints mentioned in each comment. The most prevalent issue for both groups is writing style, which is related to word choice, tone, and readability. Surprisingly, complaints like “contains things never mentioned in the paper” exist for both benign and adversarial abstracts. To investigate whether these complaints are due to LLM hallucinations during the benign paraphrasing of control abstracts rather than the original contents uploaded to arXiv, we manually examined the control and original abstracts related to each complaint. We find that all types of complaints have at least one paper which we believe the complaint applies to the original abstract.
Lastly, we investigated whether the review comments for the adversarial abstracts mentioned issues that are directly related to the modification operations proposed in this work. In the experimental group, we noticed that 6% of the responses are related to IncludeThemes, 16% of the reponses are related to InsertKeywords, and 82% of the responses are not directly related to either operations. Some responses may be related to both operations, so the percentages do not sum to 100%. Finally, none of the participants identified malicious intent in any of the abstracts. Although it is not surprising that no participant identified malicious intent given the use of deception, we believe this finding remains significant, because real-world peer reviewers likewise are not explicitly instructed to look for signs of collusion.
@@ 6 Discussion
In this work, we identify the vulnerability of current text-based reviewer-assignment systems in AI/ML venues to manipulation by collusion rings. This contradicts common perceptions that text-based automatic reviewer matching is resistant to such tactics. Our work suggests that a simple and practical attack procedure can effectively manipulate paper-reviewer similarities and hence manipulate automated reviewer assignment at many ML/AI conferences. Our attacks have a high success rate in matching to a targeted colluding reviewer. In the human subject experiment testing for attack detectability, no participant detected malicious intent for the manipulated abstracts. While there were complaints about coherence and consistency, benign abstracts edited by LLMs elicited similar complaints, suggesting plausible deniability.
This work raises awareness about the vulnerabilities of the current reviewer matching systems, and we are helping to address the these issues by informing the concerned conference management platform. To guard against attacks such as those described in this paper, platform and algorithm changes based on our findings and recommendations have been implemented platform-wide and used at a major AI/ML conference. In total, we underscore the need for enhanced robustness in reviewer assignment algorithms to protect the integrity of the peer review process.
@@ 6.1 Recommendations for mitigation
  • Increase reviewer profile requirements: Our results indicate that requiring reviewers to have more extensive publication archives reduces the effectiveness of manipulation. This makes it more difficult for colluding reviewers to align their profiles strategically with targeted abstracts.
@@
  • Increase reviewer awareness: Educating reviewers about the possibility of manipulations can prompt them to be more vigilant when evaluating abstracts and papers.
  • Introduce randomness in assignments: No component of automated assignment methods is entirely immune to manipulation. This highlights the importance of adopting a broader mitigation strategy such as introducing a degree of randomness into the assignment process [JZL+20]. Such randomness provably prevents adversaries from reliably influencing assignments.
  • Consider robustness of similarity scores: Developers of similarity algorithms should pay attention to impart some robustness so the similarity scores are less susceptible to adversarial manipulations to the abstract.
@@ 6.2 Limitations
While our study provides valuable insights, it has several limitations that offer avenues for future research. Our analysis considers scenarios where authors collude with a single reviewer within a panel of 3-6 reviewers. Expanding this to cases where multiple reviewers colluding with the author simultaneously could provide a more comprehensive understanding of the system’s vulnerabilities. We also did not factor in common constraints such as reviewer and paper load limits, conflicts of interest, or geographical considerations. Incorporating these constraints could affect the attack’s efficacy and the generalizability of some of our findings. Additionally, although the fully automatic mode enables large-scale experimentation, the abstract modifications may not fully reflect the operations of real colluders due to the absence of human supervision. Finally, it would be useful to evaluate the attack success rates in other venues.
In our human subject experiment, the adversarially modified abstract tested may be more detectable than it would be for the adversarial abstracts evaluated in all other experiments, due to the higher attack budget that was erroneously tuned for the human subject experiment. In addition, in a laboratory experiment like ours, we could not fully replicate the complexities of real-world reviewer behavior. Factors such as varying levels of expertise, bias, and attention could influence outcomes in practical settings. We also do not consider how different reviewers could influence each other.
@@ Acknowledgments
Ethical considerations
The key stakeholders of this research are ML/AI conference venues, submission authors, and its research community. As other Computer Science (CS) fields grow in size, our findings may be more widely applicable. The main risk of this work is the disclosure of a serious vulnerability in reviewer matching systems, which could allow unfair advantages and harm publication quality. To mitigate this risk, we had informed the concerned conference management platform before submission and worked with them to address these vulnerabilities using the defenses we propose. These safeguards are now in place and are being used.
Collusions and attacks described in this paper may already be happening in practice, yet most researchers may not be aware of these vulnerabilities. To promote transparency, we make our attack algorithm and adversarial abstract examples publicly available. By raising awareness of these vulnerabilities, we aim to empower non-colluding reviewers to be more vigilant. Furthermore, attacks described in this work are difficult to detect and often come with plausible deniability. Given the challenge of proving intent based on abstract content or profile changes, proactive mitigation is crucial rather than relying solely on post hoc detection and enforcement.
Finally, while this paper finds that colluders can manipulate text similarity to get high rankings with high success, it is important to note that the results do not imply any high ranking reviewer-paper pair in past data are actually colluders. Any successful attack pertains to the abstract and/or reviewer profile modified in this experiment and not to the original abstract or profile in the conference. Furthermore, the success rates are at an aggregate level, and nothing in the paper (or in any of its artifacts) should be explicitly construed to suggest collusion or a reviewing relationship.
Human Subject Experiments
This experiment was approved by an independent Institutional Review Board at Carnegie Mellon University (Federal Wide Assurance No: FWA00004206, IRB Registration No: IRB00000603). To simulate real reviewer behavior, we used deception to conceal the study’s purpose, as disclosure is likely to cause participants to act differently and be hyper-aware. All participants were debriefed and informed why deception was necessary. Participants could opt out at any time, including after debriefing, and only one participant opted-out (after debriefing). To ensure participant privacy, we only share anonymized, aggregated statistics of the collected data.
Datasets
We considered the risks of releasing reviewer and author names from the NeurIPS 2022 and 2023 datasets, given the adversarial context, but believe they are negligible. Again, we make no claims that can be construed to suggest collusion or reviewing relationships. Furthermore, all data we release is already public, making the added risk low.
Abstracts
We clearly label all modified abstracts to avoid damages to the original paper and authors. Since the original abstracts are available on OpenReview or arXiv, the risk of reputational harm is minimal. Finally, the Creative Commons Public Domain Dedication (CC0 1.0) licenses of paper metadata on OpenReview and arXiv allow us to modify and distribute the abstracts (license links: OpenReview, arXiv).
All in all, we believe the benefits of this research outweigh the risks. We hope our findings and recommendations can help the research community be more vigilant and robust to such malicious activities in scientific peer review.
Open Science
Our research artifacts listed below are available on Zenodo at https://doi.org/10.5281/zenodo.15588237.
Adversarial Examples
All adversarial abstracts evaluated in the results section are released. This includes 25 human-in-the-loop adversarial abstracts from Section 5.2; 600 fully automatic adversarial abstracts from Section 5.3, with 300 for natural ranking of 101 and 100 each for 20, 501, and 1001; 400 fully automatic adversarial abstracts from Section 5.4, with 100 each for reviewer archive lengths of 1, 2, 5, and 10; and 200 fully automatic adversarial abstracts from Section 5.5, with 100 each for average and maximum aggregation methods. The adversarial abstracts used in Section 5.6 are the same as those in Section 5.3. Additionally, all 914 fully automatic adversarial abstracts corresponding to the arXiv papers used in the human subject experiment (Section 5.7) are released.
Datasets
We release both the NeurIPS 2023 dataset and the NeurIPS 2022 dataset described in Section 5.1. We also release the dataset of 914 arXiv papers and their control abstracts (from benign LLM paraphrasing) used in Section 5.7.
Code
We release our code to execute the attacks and the notebooks/scripts used for evaluating the results in this paper. Scripts used to sample paper-reviewer pairs for evaluation (see “Evaluation samples” in Section 5.1) are also provided.
LLM Prompts
We release a total of six LLM prompts: five prompts corresponding to the IncludeThemes and InsertKeywords operations across human-in-the-loop, fully automatic, and the human study, and one prompt for paraphrasing control abstracts for the human study. Additionally, we provide four few-shot examples used in our experiments, with two examples per modification operation.
Human Subject Data
Since study participants are assigned personalized papers to review based on their past publications, the raw data of reviews, even if anonymized, can still compromise their privacy. According to IRB requirements and our promise to study participants, we only release the anonymized and aggregated statistics as presented in Section 5.7.
@@ Appendix A Adversarial Abstract Modification Algorithms
    SimilarityCheck($t_{p}$$a^{\text{adv}}_{p}$$^{\prime}$$a^{\text{adv}}_{p}$$Q_{r}^{\text{adv}}$$\delta$) queries the SPECTER model and returns true if the similarity of abstract $a^{\text{adv}}_{p}$$^{\prime}$ is higher than (or at least comparable to) the similarity of abstract $a^{\text{adv}}_{p}$, that is, $s((t_{p},{a^{\text{adv}}_{p}}^{\prime}),Q_{r}^{\text{adv}})+\delta>s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$. The $\delta$ parameter represents a small non-negative value, for cases when the new edits added in $a^{\text{adv}}_{p}$$^{\prime}$ do not have to be more similar to reviewer $r$ than the $a^{\text{adv}}_{p}$.
@@
    EarlyStoppingCheck($a^{\text{adv}}_{p}$$Q_{r}^{\text{adv}}$) returns true if the colluding reviewer $r$ (with the curated adversarial archive $Q_{r}^{\text{adv}}$) is the most-similar reviewer for the paper $p$ (with the adversarial abstract $a^{\text{adv}}_{p}$) amongst some (potentially proxy) set of reviewers. In the proposed method, abstracts are modified in a multistep process, so an early stopping check may be desirable to stop further modifications if the attack is envisaged to be successful. The use of early stopping is optional in our algorithm, since it is designed to trade off between attack effectiveness and abstract modification strength. In one of our experiments, we use the following early stopping heuristic: if the colluding reviewer is the most-similar reviewer for the colluding paper among all reviewers in the previous edition of the conference, no further abstract modifications are made. This heuristic is feasible and realistic because many major AI/ML conferences publish their reviewer pools from previous years. We validate the effectiveness of our early stopping heuristic in Section 5.2 and further investigate the use of proxy sets in Section 5.6.